OpenStudy (anonymous):

Help please. Find the mean and the range of the data set. Show your work. 55 62 61 54 68 72 59 61 70 55+62+61+54+68+72+59+61+70 = 562 55+62 = 117+61= 178+54=232+68 =300+72= 372+59=431+61=492+70=562 Our mean is 9/562 and our most common number is =61 I got 0.016014234

OpenStudy (lynfran):

to find the mean u add up all the numbers and then divide by the amount of numbers ...to find the range u take the largest number and subtract the smallest number...

OpenStudy (anonymous):

The largest number in the data set?

OpenStudy (lynfran):

yes

OpenStudy (anonymous):

70 - what?

OpenStudy (lynfran):

and ur mean shd be 562/9..... the largest #is 72

OpenStudy (anonymous):

Alright do i divide first then subtract?

OpenStudy (lynfran):

no lets do the mean first 55+62+61+54+68+72+59+61+70=562 now how many numbers are there

OpenStudy (lynfran):

how many numbers did i add to get 562?

OpenStudy (anonymous):

9

OpenStudy (lynfran):

good so we have to divide the 562 by 9 to find the mean

OpenStudy (anonymous):

562/9 = 62

OpenStudy (anonymous):

I did it backwards DX

OpenStudy (anonymous):

That's how i got a zero digit number

OpenStudy (lynfran):

62.4

OpenStudy (anonymous):

Yeah didn't think i need to include the 4 but alright.

OpenStudy (lynfran):

to find the range we take the largest # and subtract the smallest #....whats the largest number? whats the smallest #?

OpenStudy (anonymous):

72, 54

OpenStudy (lynfran):

good 72-54=?

OpenStudy (anonymous):

72-54 = 18

OpenStudy (lynfran):

good thats ur range

OpenStudy (anonymous):

Isn't all?

OpenStudy (lynfran):

if you were finding for the mode then i would have been 61...since the mode is the number that occurs the most

OpenStudy (anonymous):

Alright Thanks Lyn :)

OpenStudy (lynfran):

welcome